40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Fri Oct 4 12:33:27 2013, Masayuki, Update, SUS, Replaced the laser for Optical Lever of ETMY  OPLEV_SUM.pdfETMYoplev.png
    Reply  Fri Oct 4 14:36:44 2013, rana, Update, SUS, Replaced the laser for Optical Lever of ETMY  
    Reply  Sun Oct 6 18:37:43 2013, rana, Update, SUS, Replaced the laser for Optical Lever of ETMY  web-burnt-toast.jpg
       Reply  Mon Oct 7 10:51:18 2013, Steve, Update, SUS, Replaced the laser for Optical Lever of ETMY  ETMXoplevETMY.pngoplevSettings.png
          Reply  Tue Oct 8 08:58:16 2013, Steve, Update, SUS, Oplev power spectrums ITMoplevServoValues.pngoplevPSpectrums_darkPRM.pngBSoplevServoValues.png
             Reply  Thu Oct 10 11:30:15 2013, Steve, Update, SUS, Oplev power spectrums: PIT-YAW PitYawOLetmXY.png
             Reply  Tue Oct 15 15:25:13 2013, Steve, Update, SUS, Oplev power spectrums bandwidths BW0.01oplevPS.pngoplevPSetmXYclosed.png
                Reply  Wed Oct 16 10:32:40 2013, Steve, Update, SUS, Flowbench effect on oplev error signals FlowBenchOnOff.png
                   Reply  Thu Jul 10 15:57:00 2014, Steve, Update, SUS, flow bench is turned off 
          Reply  Thu Oct 24 17:18:44 2013, Steve, Update, SUS, ETMY oplev error signals are still bad ETMY_OPLEV_ERRORS.png
Message ID: 9202     Entry time: Fri Oct 4 12:33:27 2013     Reply to this: 9203   9206
Author: Masayuki 
Type: Update 
Category: SUS 
Subject: Replaced the laser for Optical Lever of ETMY  

[Steve, Masayuki]

We replaced the laser for optical lever of ETMY. And also we aligned the path so that beam spot hits the center for each optics. I attached the spectrum of the SUS-ETMY_OPLEV_SUM, the red curve is with old laser, and blue curve is with the new laser. We also measured without laser so as to measure the QPD dark noise (green curve). The change is significant, and seems much closer to other oplev spectrum.(Brown curve is the oplev spectrum of ITMY)

The new laser is:

Manufacture name: JDSU, Model number: 1103P, Serial number: PA892324

The injection power is 3.7 mW and the out coming power is 197 uW (measured just before the QPD). The output value of the SUS-ETMY_OPLEV_SUM is about 8500.

First we measure 2 spectrum ( including the dark noise). After that we replace the laser and align the optical lever path. We changed the post of one of the mirror (just before the QPD) because the hight was too low. Inside of the chamber is darker than before - actually we had scattering light inside the chamber before. We dumped the reflected light from QPD. And then we measured the spectrum of the oplev output. I also aligned oplev of ETMY after restoring the YARM configuration using IFO configure screen.

We don't know actually what was the problem, laser quality or the scattering light or some clipping. But the oplev seems to be better.

Steve:  Atm2 shows increased gains correction later for UGF elog 9206

Attachment 1: OPLEV_SUM.pdf  60 kB  | Hide | Hide all
OPLEV_SUM.pdf
Attachment 2: ETMYoplev.png  177 kB  Uploaded Mon Oct 7 09:55:40 2013  | Hide | Hide all
ETMYoplev.png
ELOG V3.1.3-